I am trying to reverse this array explain why it is not working and fix it thank you!
Also can you explain what ncurses.h is and its utalizations. Thanks!


#include <stdio.h>

array[10] = {1,2,3,4,5,6,7,8,9,10};
int i,j,temp;

int
main(){

    for(i = 0; i < 10; i++) {
        for (j = 9; j >= 0; j--) {
            temp = array[i];
            array[i] = array[j];
            array[j] = temp;

        }
    }





    for(i = 0; i < 10; i++) {
            printf(" %d", array[i]);
        }
    }
Answer #1

If the answer helps you please upvote, and in case you have any doubts please comment and get it clarified,.

You are using the wrong algorithm to reverse the array. We use only one loop to reverse an array.

Swap elements in the first half with elements in the second half to reverse the array.

element at index 0 = 1 swapped with element at index 9-0 = 9 = 10
element at index 1 = 2 swapped with element at index 9-1 = 8 = 9
element at index 2 = 3 swapped with element at index 9-2 = 7 = 8
element at index 3 = 4 swapped with element at index 9-3 = 6 = 7
element at index 4 = 5 swapped with element at index 9-4 = 5 = 6

Output:

Ncurses is a library in c, it is accessed using ncurses.h. It is a library used for developing GUI_like application software that runs under a terminal emulator. It is used for developing terminal applications.